# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ Abstract
Spectral partitioning is a method that can be used to compute small sparse cuts or small edge-separators in a wide variety of graph classes, by computing the second-smallest eigenvalue (and eigenvector) of the Laplacian matrix. Upper bounds on this eigenvalue for certain graph classes imply that the method obtains small edge-separators for these classes, usually with a sub-optimal dependence on the maximum degree. In this work, we show that a related method, called reweighted spectral partitioning, guarantees near-optimal sparse vertex-cuts and vertex-separators in a wide variety of graph classes. In many cases, this involves little-to-no necessary dependence on maximum degree.
We also obtain a new proof of the planar separator theorem, a strengthened eigenvalue bound for bounded-genus graphs, and a refined form of the recent Cheeger-style inequality for vertex expansion via a specialized dimension-reduction step.
@@ 1 Introduction
Partitioning a graph into smaller pieces is a fundamental type of problem that arises in a variety of fields. Applications of graph partitioning include divide and conquer algorithms, dynamic programming algorithms [RH01], graph drawing [BPP22], geometrically embedding graphs [SSX25], FPGA compilation [CCW21], image processing, VLSI design, compute parallelization, bioinformatics, simulation, road networks, social networks [BMS+16], graph processing [OFS17], and even air traffic control methods [FGBJ17].
Many of these applications have similar ways of measuring partitioning quality, giving rise to some notable well-studied objectives, such as various forms of “expansion” and “separators”, each of which is a class of objectives with a variety of theoretical applications. In their most fundamental form, these objective types are studied in the sense of a single “cut”, partitioning the graph into two parts.
Fix a graph $G=(V,E)$ with $n$ vertices. For a subset $A\subset V$ with $|A|\leq n/2$, let $\delta(A)$ denote the set of edges with one endpoint in $A$ and one endpoint in $V\setminus A$. Likewise, let $\partial A$ denote the set of vertices in $V\setminus A$ adjacent to some vertex in $A$. The edge expansion of $A$ is the ratio $\frac{|\delta(A)|}{|A|}$. The vertex expansion of $A$ is the ratio $\frac{|\partial(A)|}{|A|}$. Due to this form, $A$ is often called a ratio cut in both cases, and a bounded-ratio cut if only an upper bound on the expansion is known. The edge or vertex expansion of $G$ is defined as the minimum edge or vertex expansion over all possible choices of $A$. The edge expansion of $G$ is denoted $\phi(G)$, and the vertex expansion of $G$ is denoted $\psi(G)$. Edge and vertex expansion are both NP-hard to compute exactly [Kai04] (see Appendix B for vertex expansion). Moreover, for a graph with vertex expansion $\psi$ and maximum degree $\Delta$, it is hard to find a subset with vertex expansion less than $C\sqrt{\psi\log\Delta}$ for an absolute constant $C$, assuming the so-called “small-set expansion hypothesis” [LRV13], which was suggested by Raghavendra and Steurer [RS10].
@@
{TAB}
(5,3pt,3pt)[5pt]—l—c—c—c——c—c—c—c—c—c—c— Graph class Separator Size Time Refs.
Planar $\operatorname{\mathcal{O}}(\sqrt{n})$ $\operatorname{\mathcal{O}}(n)$ [Ung51, LT79]
[…]
$K_{h}$-minor-free $\operatorname{\mathcal{O}}(h\sqrt{n})$ $\operatorname{\mathcal{O}}(n^{1+\varepsilon})$ [^2] Assumes $h$ is a constant, and uses a tower function in terms of $h$. Note that some more tractable algorithms are known with weaker guarantees [WN11]. [KR10]
RIG over $K_{h}$-minor-free $\operatorname{\mathcal{O}}\!\bigl{(}h^{3}\sqrt{m\log h}\bigr{)}$ $\text{poly}(n)$ [^3] Solves a sequence of $O(n)$ linear programs, each with approximately $n^{2}$ variables and $n^{3}$ (sparse) constraints (exact number of each depends on formulation, but is always polynomial). [Lee17]
$k$-ply neighbourhood system in $\mathbb{R}^{d}$ $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$ $O(nd^{2})$ [^4] Assumes an embedding of the graph in $\mathbb{R}^{d}$ is given. [MTTV97]
$k$-nearest-neighbour graph in $\mathbb{R}^{d}$ $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$ $O(nd^{2})$ [^4] Assumes an embedding of the graph in $\mathbb{R}^{d}$ is given. [MTTV97]
Table 1: Best-known sizes of 2/3 2/32 / 3-vertex-separators for several graph classes, as well as the time required to compute the separator. Many of these results require additional information besides the graph to compute the separator in the stated time complexity (see the footnotes). $n$ is always the number of vertices, and $m$ is always the number of edges.
These two types of objectives are intimately connected. An edge or vertex-separator $S$ itself forms either $S=\delta(A)$ or $S=\partial(A)$ (respectively) for some set $|A|\leq n/2$, so a graph with an edge/vertex-separator $S$ has edge/vertex-expansion at most $O\left(|S|/n\right)$. A converse form of this is also often true. In particular, it can be shown by a standard recursive argument that if every induced subgraph $H\subset G$ has edge/vertex expansion at most $\kappa$, then $G$ has an edge/vertex separator of size $O(\kappa n)$ (see e.g. [ST96, ST07, Lemma A.1]).
1.1 Algorithms for General Graphs
Although all types of expansion and separators we have mentioned are NP-hard to compute in general, a number of different algorithms with guarantees are known. These algorithms use no special information about the graph, but still have guarantees for many graph classes.
1.1.1 Spectral Partitioning

### v4

# Reweighted Spectral Partitioning Works: A Simple Algorithm for Vertex Separators in Special Graph Classes

@@ Abstract
We establish that a simple polynomial-time algorithm that we call reweighted spectral partitioning obtains small $\frac{2}{3}$-balanced vertex-separators for a number of graph classes, including $\operatorname{\mathcal{O}}(\sqrt{n})$-sized separators for planar graphs, $\operatorname{\mathcal{O}}(\min\{(\log g)^{2},\log\Delta\}\cdot\sqrt{gn})$-sized separators for genus-$g$ graphs of maximum degree $\Delta$, and $\operatorname{\mathcal{O}}(\min\{\log h,\sqrt{\log\Delta}\}(h\log h\log\log h)\sqrt{n})$-sized separators for $K_{h}$-minor-free graphs of maximum degree $\Delta$.
To accomplish this, we first obtain a refined form of a Cheeger-style inequality relating the vertex expansion of a graph and the solution to a semidefinite program defined over the graph. Then, to obtain the guarantees for specific graph classes, we derive direct bounds on the value of the semidefinite program.
We also obtain several other results of independent interest, including an improved separator theorem for the intersection graphs of $d$-dimensional balls with bounded ply, a new bound on the Fiedler value of genus-$g$ graphs, and a new “spectral” proof of the planar separator theorem.
@@ 1 Introduction
In this work, we study simple algorithms for computing small balanced vertex separators.
Let $G=(V,E)$ be a graph with $n$ vertices and $m$ edges. For a parameter $\alpha\in(\frac{1}{2},1)$, a subset of the vertices $S\subset V$ is called an $\alpha$-balanced vertex-separator of $G$ if every connected component of the induced subgraph of $G$ on $V\setminus S$ has at most $\alpha\cdot n$ vertices. An $\alpha$-balanced edge-separator of $G$ is a subset of edges meeting the same criteria. For simplicity, we will sometimes use the term separator to refer to a $\frac{2}{3}$-balanced vertex-separator.
Small separators are useful for a vast number of applications, including divide and conquer algorithms and dynamic programming [RH01]. Moreover, small separators are known to exist for a large number of common graph classes (see Appendix A for definitions of various graph classes). Most notably, the planar separator theorem states that a planar graph with $n$ vertices has a separator consisting of $\operatorname{\mathcal{O}}(\sqrt{n})$ vertices [Ung51, LT79]. Such a separator can also be found in linear time. Similar existential results are also known for other graph classes, such as (oriented) genus-$g$ graphs always having separators consisting of $\operatorname{\mathcal{O}}(\sqrt{g\,n})$ vertices [GHT84], and $K_{h}$-minor-free graphs always having separators consisting of $\operatorname{\mathcal{O}}(h\sqrt{n})$ vertices [KR10]. Both of these example bounds are also tight. Unfortunately, while all of these existential results also imply algorithms, these algorithms also require time exponential in the class parameter ($g$ or $h$).
Since the implicit algorithms for producing these separators are slow, it is natural to consider more general approximation algorithms. However, approximating the minimum-size $\alpha$-balanced vertex-separator of a graph is NP-hard even for graphs of maximum degree $3$, and even up to an $\operatorname{\mathcal{O}}(n^{\frac{1}{2}-\epsilon})$ additive approximation [BJ92] (for any $\epsilon>0$). Fortunately, pseudo-approximations that relax the balance $\alpha$ are just as useful for most applications: For a graph admitting a $2/3$-balanced vertex-separator of size $k$, an algorithm of Feige, Hajiaghayi, and Lee [FHL05, FHL08] can be used to produce a $3/4$-balanced separator of size $\operatorname{\mathcal{O}}(k\cdot\sqrt{\log n})$ in polynomial time. In fact, they can obtain a separator of size $\operatorname{\mathcal{O}}(k\cdot\sqrt{\log k})$.
We study a slightly different approach. Rather than aim for general (pseudo-)approximation algorithms or specialized per-class algorithms, we will study a polynomial-time algorithm for producing $\frac{2}{3}$-balanced vertex-separators in an arbitrary graph. We call this algorithm reweighted spectral partitioning due to its numerous similarities with the popular “spectral-partitioning” algorithm [ST96, ST07]. This algorithm is sufficiently simple that we have implemented it, [^1] We have not implemented an optional partition oracle used by the algorithm. and examples of separators that it produces can be found in Figure 1.
Figure 1: Examples of vertex separators produced by reweighted spectral partitioning: A $100$-vertex grid graph (left), a $50$-vertex random planar graph (centre), and $50$-vertex 22-dimensional 55-nearest-neighbour graph (right). The separator vertices are the large red ones.
Compared to the pseudo-approximations mentioned above, reweighted spectral partitioning has relatively weak guarantees for general graphs. However, the primary goal of this paper is to show that it admits quite strong guarantees for a number of specific graph classes, including planar graphs and bounded-genus graphs. In fact, we will show that the separators it produces nearly match the per-class separator sizes for a number of important graph classes.
1.1 The Method of Reweighted Spectral Partitioning
The reweighted spectral partitioning algorithm that we consider is a rounding algorithm for a semidefinite program. In particular, it is the combination of algorithms implicit in the proofs of Kwok, Lau, and Tung [KLT22] and Biswal, Lee, and Rao [BLR10]. Critically, it makes use of the following optimization problem parametrized by a value $d\in\mathbb{N}$:
@@ 1.1 The Method of Reweighted Spectral Partitioning
Several previous works have studied this problem for $d=n$ or $d=1$ [BDX04, Roc05, OTZ22, JPV22, KLT22]. In particular, when $d=n$, this is a semidefinite program and it is known to satisfy Slater’s condition, so it can be approximately solved in polynomial time with general methods [JKL+20].
The reweighted spectral partitioning algorithm we consider then consists of four high-level steps for a graph $G$:
  1. 1.
    Approximately solve the semidefinite program $\gamma^{(n)}(G)$, obtaining variables $f(v)\in\mathbb{R}^{n}$ and $y(v)\in\mathbb{R}_{\geq 0}$ for each vertex $v\in V$.
  2. 2.
    Dimension-reduce the $f(v)$ variables to get one-dimensional values $f^{\prime}(v)\in\mathbb{R}$. Several methods/samples for dimension-reduction are applied, and the best result is used. One of the methods takes, as a parameter, an optional oracle that partitions the vertices of the graph based on a vertex-weighting.
  3. 3.
    Sort the vertices by their $f^{\prime}(v)$ values to get an ordering $v_{1},\dots,v_{n}$, and consider partitions of the form $A_{k}=\{v_{1},\dots,v_{k}\},B_{k}=\{v_{k+1},\dots,v_{n}\}$. For each crossing edge in $A_{k}\times B_{k}\cap E$, select its incident vertex $v$ maximizing $y(v)$, and let $S_{k}$ be the set of selected vertices. Return $(A,S,B):=(A_{k},S_{k},B_{k})$ minimizing $\frac{|S_{k}|}{\min\{|A_{k}|,|B_{k}|\}}$.
  4. 4.
    Repeat on the subgraph of $G$ induced by larger of the sets $A$ and $B$ until the union of all the computed sets $S$ forms a balanced separator.
Importantly, the algorithm requires no information about the graph other than its vertices and edges. The optional partitioning oracle could make use of such information (if desired), but the useful partitioning oracles we will suggest do not. This makes the algorithm suitable for very general applications.
1.2 Main Results
We will show that the reweighted spectral partitioning algorithm expressed in the previous subsection produces provably good results in a number of different senses. Some of these results will make use of a specialized random partition oracle as a subroutine, while others will not.
We will use the notation $a\lesssim b$ to mean $a=\operatorname{\mathcal{O}}(b)$.
1.2.1 Refined Cheeger-Style Inequality
The first main result is an inequality relating $\gamma^{(n)}$ and $\gamma^{(1)}$ with several important quantities that we will now define. The most important of these is a quantity related to vertex separators:
Definition 1.1.
For a subset $A\subset V$ with $|A|\leq n/2$, let $\partial A$ denote the set of vertices in $V\setminus A$ adjacent to some vertex in $A$. The vertex expansion of a set $A$ is the ratio $\psi(A):=\frac{|\partial A|}{|A|}$. The vertex expansion of a graph $G$ is the minimum vertex expansion over all possible sets $|A|\leq\frac{n}{2}$, and it is denoted as $\psi(G)$.
Computing vertex expansion is known to be NP-hard (see Appendix B for details), so it is usually studied alongside approximation algorithms. Vertex expansion is intimately related to balanced vertex-separators. Notably, approximations of vertex-expansion imply pseudo-approximations of balanced vertex-separators [FHL05, FHL08, LR99]. In our case, a simpler fact is more relevant: Given a graph $G$ with $n$ vertices, if every induced subgraph $H$ of $G$ on $n^{\prime}$ vertices has vertex expansion at most $\frac{\kappa}{\sqrt{n^{\prime}}}$, then $G$ admits a $\frac{2}{3}$-balanced vertex separator of size $2\kappa\sqrt{n}$. This separator can be obtained by repeatedly finding sets $A$ of vertices with small vertex expansion, and then deleting $\partial(A)$ and $A$ from the graph. The final separator is the union of all sets $\partial A$ from this process. This reduction is standard for a number of different formulations of expansion and separators (e.g., see [ST07, Lemma A.1]).
The second quantity relates to our optional partition oracle used for dimension-reduction. Intuitively, the following quantity arises when we aim to randomly partition a metric while trying to avoid cutting large metric balls (in expectation):
Definition 1.2.
Let $(X,d)$ be a metric space. The (weak) modulus of padded decomposability of $(X,d)$ is defined as the smallest value of $\alpha$ so that for every $\Delta>0$, there exists a distribution $\mu$ of partitions of $X$ with parts of diameter at most $\Delta$, and
where $B(x,r)$ denotes the metric ball of radius $r$ centred at $x$. We denote this quantity as $\alpha(X,d)$.
Definition 1.3.
For a graph $G=(V,E)$, the extremal vertex modulus of padded decomposability on $G$ is defined as the maximum value of $\alpha(V,d_{\omega})$ for vertex-weight functions $\omega:V\to\mathbb{R}_{\geq 0}$ inducing shortest-path metrics $d_{\omega}$ on $G$. We denote this quantity as $\alpha(G)$.
A similar definition exists for edge-weighted shortest-path metrics, but we will only use this one. $\alpha(G)$ is known to be bounded for a number of graph classes. In particular, for $K_{h}$-minor-free graphs, $\alpha(G)\in\operatorname{\mathcal{O}}(\log h)$ [CF25]. Note that genus-$g$ graphs are $K_{\Theta(\sqrt{g})}$-minor-free.
Finally, the result we will prove is:
Theorem 1.4 (Refined Cheeger-Style Inequality for Vertex Expansion).
For a graph $G$ with $n$ vertices and maximum degree $\Delta$,
$\frac{\psi(G)^{2}}{\min\{\log\Delta,[\alpha(G)]^{2}\}}\lesssim\frac{\gamma^{(1)}(G)}{\min\{\log\Delta,[\alpha(G)]^{2}\}}\lesssim\gamma^{(n)}(G)\lesssim\gamma^{(1)}(G)\lesssim\psi(G).$
Moreover, given an oracle that produces an $\alpha$-padded partition for any vertex-weighting of $G$ in polynomial time, the reweighted spectral partitioning algorithm can compute in (Monte Carlo randomized) polynomial time a set $S$ with $|S|\leq\frac{n}{2}$ and vertex expansion $\psi(S)\lesssim\sqrt{\gamma^{(n)}(G)\cdot\min\{\log\Delta,\alpha^{2}\}}$.
This is referred to as a “Cheeger-style” inequality since the upper bounds on $\psi(G)$ are given in terms of a square root of an expression involving either $\gamma^{(1)}(G)$ or $\gamma^{(n)}(G)$, while the lower bounds on $\psi(G)$ have no such square root operation. This is similar to the well-known Cheeger inequality, which relates some similar quantities [HLW06, Theorem 4.11].
The proof of this inequality will directly use the algorithm of reweighted spectral partitioning (more specifically, Algorithm 5 from Section 3). In fact, this result is a refinement of a result from a long sequence of Cheeger-style inequalities for vertex-expansion [Roc05, OTZ22, JPV22, KLT22]. In particular, the specific algorithm we use is a refinement of the one implicit in the proofs of Kwok, Lau, and Tung [KLT22], incorporating a new dimension-reduction step based on a technique of Biswal, Lee, and Rao [BLR10] that uses random partitions. Their algorithm is not the only one that could form the basis of ours, but it has the benefit of being quite simple.
1.2.2 Vertex Separators for Special Graph Classes
The bulk of this work is dedicated to showing that reweighted spectral partitoning can be used to recover balanced vertex separators for a number of special graph classes that have nearly-optimal size for that class. All of these proofs will make use of part of the Cheeger-style inequality: By directly bounding $\gamma^{(d)}$ for a particular class (and some dimension $d$), we will obtain guarantees for the performance of reweighted spectral partitioning on that class by applying Theorem 1.4. We present several such results.
Theorem 1.5 (Bounds on $\gamma^{(d)}$).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. Then the following conditional bounds hold:
  • If $G$ is planar, then $\gamma^{(1)}(G)\lesssim\frac{1}{n}$.
  • If $G$ has (oriented) genus $g$, then $\gamma^{(1)}(G)\lesssim\frac{g\cdot\min\{(\log g)^{2},\log\Delta\}}{n}$.
  • If $G$ is $K_{h}$-minor-free, then $\gamma^{(1)}(G)\lesssim\frac{(h\log h\log\log h)^{2}}{n}$.
  • If $G$ is a $d$-dimensional $k$-ply ball-intersection graph [^2] Sometimes called a neighbourhood-system. in $\mathbb{R}^{d}$, then $\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}$.
  • If $G$ is a $d$-dimensional $k$-nearest-neighbour graph in $\mathbb{R}^{d}$, then $\gamma^{(d+1)}(G)\lesssim\left(\frac{k}{n}\right)^{\frac{2}{d}}$.
All of these classes are hereditary (closed under induced subgraphs), so these bounds can be combined with Theorem 1.4 in several different ways to obtain results for balanced vertex-separators. The first is a purely existential construction that uses the bound $\psi(G)^{2}\lesssim\gamma^{(d)}(G)\cdot\min\{\log\Delta,[\alpha(G)]^{2}\}$:
Corollary 1.6 (Existential Separator Bounds).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. Then the following conditional bounds hold:
  • If $G$ is planar, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{n})$.
  • If $G$ has (oriented) genus $g$, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{\log g,\sqrt{\log\Delta}\}\sqrt{gn})$.
  • If $G$ is $K_{h}$-minor-free, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}((h\log h\log\log h)\sqrt{n})$.
  • If $G$ is a $k$-ply ball-intersection graph in $\mathbb{R}^{d}$, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d})$.
  • If $G$ is a $k$-nearest-neighbour graph in $\mathbb{R}^{d}$, then $G$ admits a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\min\{d,\log\Delta\}}\cdot k^{1/d}n^{1-1/d})$.
Take note of the fact that these bounds do not result from running reweighted spectral partitioning, and are not inherently constructive. Moreover, the first three of these five bounds are not new, and the bounds for genus-$g$ and $K_{h}$-minor-free graphs are also weaker than the known existential bounds of $\operatorname{\mathcal{O}}(\sqrt{gn})$ [GHT84] and $\operatorname{\mathcal{O}}(h\sqrt{n})$ [KR10], respectively. However, the final two of these bounds for the geometric classes are brand new separator bounds, improving over the previously-known bound of $\operatorname{\mathcal{O}}(dk^{1/d}n^{1-1/d})$ [MTTV97]. While separators of this size would not result from reweighted spectral partitioning, we will discuss their algorithmic aspects in Section 2.3.2.
The second set of bounds we obtain for separators are those that are achieved by reweighted spectral partitioning when supplied with the polynomial-time random padded partition oracle of Conroy and Filtser [CF25]:
Corollary 1.7 (Separator Bounds via Reweighted Spectral Partitioning with Padded Partitions).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. When provided with a known polynomial-time algorithm to (randomly) partition the vertex weights, reweighted spectral partitioning can be used to compute the following sizes of separators in (Monte Carlo randomized) polynomial time:
  • If $G$ is planar, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{n})$.
  • If $G$ has (oriented) genus $g$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{(\log g)^{2},\log\Delta\}\sqrt{gn})$.
  • If $G$ is $K_{h}$-minor-free, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{\log h,\sqrt{\log\Delta}\}(h\log h\log\log h)\sqrt{n})$.
The final set of bounds we obtain for separators are those that are achieved by reweighted spectral partitioning when no partitioning oracle is provided:
Corollary 1.8 (Separator Bounds via Reweighted Spectral Partitioning with No Partition Oracle).
Let $G$ be a graph with $n$ vertices and maximum degree $\Delta$. Then reweighted spectral partitioning can be used to compute the following sizes of separators in (Monte Carlo randomized) polynomial time:
  • If $G$ is planar, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{n\log\Delta})$.
  • If $G$ has (oriented) genus $g$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\min\{\log g\cdot\sqrt{\log\Delta},\log\Delta\}\cdot\sqrt{gn})$.
  • If $G$ is $K_{h}$-minor-free, then it produces a $\frac{2}{3}$-balanced vertex-separator of size
    $\operatorname{\mathcal{O}}(\sqrt{\log\Delta}\cdot(h\log h\log\log h)\cdot\sqrt{n})$.
  • If $G$ is a $k$-ply ball-intersection graph in $\mathbb{R}^{d}$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\log\Delta}\cdot k^{1/d}n^{1-1/d})$.
  • If $G$ is a $k$-nearest-neighbour graph in $\mathbb{R}^{d}$, then it produces a $\frac{2}{3}$-balanced vertex-separator of size $\operatorname{\mathcal{O}}(\sqrt{\log\Delta}\cdot k^{1/d}n^{1-1/d})\leq\operatorname{\mathcal{O}}(\sqrt{d+\log k}\cdot k^{1/d}n^{1-1/d})$.
Not providing an oracle has the benefit of keeping the algorithm incredibly simple (see Section 3), but the guarantees on the size of the resulting separators are weaker.
1.2.3 Comparison to Other Algorithms
The main benefit of reweighted spectral partitioning is that it is a simple polynomial-time algorithm that provides near-optimal separators for a number of graph classes, and requires nothing but the graph itself as input. However, despite its simplicity, it remains comparable with other known polynomial-time algorithms that do not require anything but the graph itself, and in some cases it produces improved results. We provide comparisons to other work in Table 1.
Table 1: Comparison of our performance guarantees for reweighted spectral partitioning with existing (Monte Carlo randomized) polynomial-time methods. The best bound is always bolded (multiple are bolded if the best bound depends on parameters). Note that one of the previous works’ bounds for genus-$g$ graphs requires a fix given in Section 5.3.
2 Technical Overview, Related Work, and Additional Results
In this section, we will briefly outline some related work in spectral graph theory, and then discuss our techniques for obtaining bounds. In particular, we will use two different families of techniques for obtaining bounds, and we will discuss each one separately.
2.1 Fiedler Bounds and the Cheeger Inequality